OpenStudy (anonymous):

Please help!!!! Which graph represents the solutions to the inequality |4x - 8| > 8

OpenStudy (anonymous):

@amistre64 @thomaster @undeadknight26 @ParthKohli @SolomonZelman @ganeshie8 @Hero @Luigi0210 @Loser66 @triciaal

OpenStudy (amistre64):

the one that cuts |4x-8| off at y=8

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

@amistre64 I don't understand

OpenStudy (amistre64):

oh, well it looks like you narrowed it down some, can you tell me why you want to pick B?

OpenStudy (anonymous):

well its not going to be a closed circle right because its greater than and not greater than or equal to

OpenStudy (amistre64):

good, what else do we know .... that will help us out? we know that 0<8 so try to solve 4x-8 = 0 to see what NOT to include

OpenStudy (anonymous):

x=2?

OpenStudy (amistre64):

yes x=2 makes it 0, so we do NOT want the section that has 2 in it .... B fits the bill perfectly.

OpenStudy (anonymous):

o thank you so much

OpenStudy (amistre64):

well, there is no color legend, im assuming blue are solutions, and yellow are not?

OpenStudy (amistre64):

otherwise its hard to tell if we pick A or B .... depends on what color defines the solution area.

OpenStudy (anonymous):

yes the blue are the solutions

OpenStudy (amistre64):

yay!! lol

OpenStudy (triciaal):

first let 4x -8 = 0 x=2 when x = 0 y = -8 but use the absolute value = 8 draw the line using points (08) and (2,0) shade the section to the right ask for greater than 0

OpenStudy (triciaal):

sorry i had 4x -8 >0 it shoud be>8 to give the point (4,0) won't change the answer still B

OpenStudy (anonymous):

thank you could you help me with another

OpenStudy (triciaal):

post

OpenStudy (triciaal):

that's the same one

OpenStudy (anonymous):

no its a different one

OpenStudy (triciaal):

now it is 2x -4 <=14

OpenStudy (anonymous):

yes

OpenStudy (triciaal):

use the same process 2x - 4 + 4 = 14 + 4 x = 9 boundary

OpenStudy (triciaal):

when x = -9 y = 22

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

C include the points shaded circle when x = -5 solution -5 to x = 9 include

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

wait so which one would it be

OpenStudy (triciaal):

sorry B reason correct

OpenStudy (triciaal):

so very sorry i am getting confused when we have absolute we are taking the positive value

OpenStudy (triciaal):

last one was easy i drew the graph and then read the numbers

OpenStudy (anonymous):

o ok thank you so much

OpenStudy (triciaal):

let me do the graph to confirm give me a few minutes

OpenStudy (triciaal):

-5 to 9 included

OpenStudy (triciaal):

already showed when x = -9 value =22 so not included limit is -5